# HET/HPF observations of Helium in warm, hot, and ultra-hot Jupiters

### v1

@@ Abstract
The near-infrared helium triplet line is a powerful tool for studying atmospheric escape processes of close-in exoplanets, especially irradiated gas giants. Line profile fitting provides direct insight into the mechanisms driving atmospheric mass loss of close-in, Jupiter-sized planets. We present high-resolution transmission spectroscopy results for the helium triplet line of sixteen gas giants (). These observations are part of an extensive helium survey conducted using the Habitable Zone Planet Finder spectrograph on the 10 m Hobby-Eberly Telescope. For the first time, we provide constraints on the helium line for HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. Additionally, we are able to confirm previous robust or tentative detections for HD189733 b, HD209458 b, WASP-52 b, WASP-69 b, and WASP-76 b, and non-detections for HAT-P-3 b, WASP-11 b, WASP-80 b, WASP-127 b, and WASP-177 b. We do not confirm the previous helium narrow-band detection in HAT-P-26 b using high-resolution observations. To identify trends within the population of warm, hot, and ultra-hot Jupiters, we combined our results with available helium studies from the literature. As predicted by theory, we find that warm Jupiters with helium detections orbit K-type stars. However, the helium detections in hot Jupiters are found in low-density planets orbiting F-type stars. There are no helium detections in ultra-hot Jupiters. We compiled a list of 46 irradiated gas giants, but more helium studies are needed to increase the sample and improve our understanding of atmospheric mass loss through helium observations.
@@ 1 Introduction
Gas giant exoplanets, particularly those on short‑period orbits, are the primary laboratories for atmospheric characterization because their large radii and extended atmospheres produce deep and readily detectable transit signals (S. Seager & D. Deming, 2010). Due to their proximity to the host star, they can reach high equilibrium temperatures (), ranging from warm Jupiters (WJs, 1000 K) through hot Jupiters (HJs, 1000–2000 K) to ultra‑hot Jupiters (UHJs, 2150 K, V. Parmentier et al., 2018; M. Stangret et al., 2022). During a transit, starlight filters through the planetary atmosphere, enabling transmission spectroscopy with high-resolution spectrographs, which probes the composition of the upper atmospheric layers. The combination of large scale heights, high irradiation, and short orbital periods makes warm, hot, and ultra-hot Jupiters particularly favorable for atmospheric studies (I. J. M. Crossfield, 2015; N. Madhusudhan, 2019). These planets have provided the first detections of exoplanetary atomic species, and escaping material (D. Charbonneau et al., 2002; A. Vidal-Madjar et al., 2003).
@@
The atmospheric escape of close-in planets is primarily hydrodynamic, driven by the thermal energy deposited into the exoplanet’s atmosphere (R. T. Pierrehumbert, 2010; T. T. Koskinen et al., 2014). Photo-evaporation is the typical hydrodynamic escape process and is driven by stellar irradiation in the most energetic wavelengths, i.e., the X-ray and extreme ultraviolet (XUV111We define XUV as wavelengths up to 504 Å since this range is relevant to helium studies due to the ionization wavelength cutoff of the He i. We note that other works may define the XUV range different, e.g. up to the H ionization cutoff at 912Å.) range (J. E. Owen & A. P. Jackson, 2012). This energy input can inflate the upper atmosphere carrying light elements, such as hydrogen and helium, to higher altitudes. Additionally, the outer layers of the atmosphere of extremely close-in planets can extend beyond the Roche lobe, resulting in a high atmospheric mass loss rate (J. E. Owen, 2019). Roche lobe overflow escape is independent of stellar flux and can act alongside other evaporation mechanisms (A. Lecavelier des Etangs et al., 2004). In extreme cases, these extended, escaping atmospheres can result in the formation of comet‑like tails that interact with the stellar wind, detectable via atmospheric escape tracers, such as the He i triplet line (e.g., Z. Zhang et al., 2023; M. Gully-Santiago et al., 2024). Over time, these mass‑loss processes can sculpt planetary demographics, contributing to observed features such as the sub‑Jovian desert, and providing empirical constraints for models of photo-evaporation (J. E. Owen & A. P. Jackson, 2012; J. E. Owen & D. Lai, 2018).
@@
The He i triplet at 10833 Å has become a key diagnostic for probing the upper atmospheres and escape processes of close‑in exoplanets, including irradiated gas giants (S. Seager & D. D. Sasselov, 2000; A. Oklopčić & C. M. Hirata, 2018; M. Lampón et al., 2021). The absorption arises from the metastable 23S state of neutral helium, which is populated when stellar XUV radiation ionizes helium atoms, and subsequent recombination fills the triplet state. However, the helium metastable state has a lifetime of 2.2 h (G. W. Drake, 1971). Thus, a constant and intense irradiation is needed to sustain detectability of the He i triplet line (J. Sanz-Forcada & A. K. Dupree, 2008). Since its first detection in WJs (J. J. Spake et al., 2018, WASP-107 b; L. Nortmann et al., 2018, WASP-69 b), helium absorption has been reported in various types of exoplanets, providing a direct tracer of atmospheric escape accessible from ground-based telescopes with high‑resolution infrared spectrographs (J. Orell-Miquel et al., 2024). Moreover, ground-based high-resolution spectroscopy observations allow to retrieve several physical parameters from line profile fitting to study the hydrodynamical escape (e.g. M. Lampón et al., 2023; J. Orell-Miquel et al., 2023). Therefore, helium spectroscopy provides a unique window into both the instantaneous escape dynamics and the long‑term sculpting of exoplanet populations.
@@
In this context, this work presents the results of an observational study aimed at characterizing atmospheric escape in gas giants. We have been carrying out a survey targeting the He i triplet line for planets with 0.5 RJup and ranging 800–2100 K around F-, G-, and K-type host stars. Figure 1 presents the planets observed by our survey. The individual analyses and discussions for HAT-P-32 b and HAT-P-67 b are presented in Z. Zhang et al. (2023), and M. Gully-Santiago et al. (2024), respectively. Here, we only included those results, and not new analysis on these two systems in our helium detections discussion.
2 Observations and data analysis
2.1 HPF spectroscopic observations
@@
We obtained spectroscopic observations of gas giant host stars with the Habitable Zone Planet Finder (HPF; S. Mahadevan et al. 2012, 2014; A. J. Metcalf et al. 2019) mounted on the 10 m Hobby-Eberly Telescope (HET; L. W. Ramsey et al. 1998; M. Shetrone et al. 2007). HPF spectrograph covers the near-infrared wavelength range from 0.81–1.38 m with a resolving power of 55 000. HPF data were reduced using the Goldilocks222https://github.com/grzeimann/Goldilocks_Documentation pipeline (J. P. Ninan et al., 2018; K. F. Kaplan et al., 2019), obtaining the science, blank sky, and calibration. To avoid saturation or scattered light from the laser frequency comb, we did not use a simultaneous calibration source in the calibration fiber.
@@
The HET telescope structure can rotate in azimuth but it is fixed at an elevation of 55 degrees. Target tracking is performed by a prime focus instrument package that moves across the focal plane to follow the target. Due to the design of the HET, we are limited to observations during “tracks” of typically 1 hour before (east track) and after (west track) target culmination, making it difficult observe a full transit, or to obtain in-transit and out-of-transit spectra on the same night. Out-of-transit phases are usually covered by observations during the nights surrounding each transit event. HPF observations were organized into visit campaigns around an exoplanet transit. We also obtained random-phase spectra, referred to as P4 observations, to monitor long-term stellar variability. P4 observations are typically used to fill the HET queue or when the conditions are not ideal for higher-priority targets.
@@
We provide a summary of the observations performed for each target analyzed in this work. We report the signal-to-noise ratio (S/N) per extracted pixel of the échelle order where the He triplet line is located.
HAT-P-3 b: We scheduled one visit, centered on the transit on 1 February 2021, during which we collected 16 spectra: 8 in-transit and 8 out-of-transit; complemented with 4 P4 spectra. We set the exposure time to 630 s, achieving a median S/N of 94.
HAT-P-12 b: We collected 21 spectra on visit 1 (31 March 2020): 5 in-transit and 16 out-of-transit; and 12 spectra on visit 2 (27 May 2025): 8 in-transit and 4 out-of-transit. We also collected 2 P4 spectra. We set the exposure time to 960 s for visit 1 and 620 s for visit 2, achieving a median S/N of 52 in both visits.
HAT-P-17 b: We collected 23 spectra on visit 1 (8 June 2020): 4 in-transit and 19 out-of-transit; and 32 spectra on visit 2 (30 July 2020): 7 in-transit and 25 out-of-transit. We also collected 8 P4 spectra. We set the exposure time to 310 s for both visits, achieving a median S/N of 78 and 117, respectively.
HAT-P-26 b: We scheduled one visit, centered on the transit on 1 February 2021, during which we collected 15 spectra: 5 in-transit and 10 out-of-transit; complemented with 2 P4 spectra. We set the exposure time to 660 s, achieving a median S/N of 89.
@@
HD118203 b: We collected 26 spectra on visit 1 (7 January 2021): 14 in-transit and 12 out-of-transit; and 29 spectra on visit 2 (19 February 2021): 18 in-transit and 11 out-of-transit. We set the exposure time to 310 s for both visits, achieving a median S/N of 325.
HD189733 b: We obtained 9, 9, and 2 in-transit spectra for the transits on 4 May 2019, 13 June 2019, and 1 August 2020, respectively. We only collected 11 spectra out-of-transit on the night of 17 April 2019, complemented with 10 P4 spectra across 2020. We set the exposure time to 310 s, achieving a typical median S/N of 400.
HD209458 b: We scheduled one visit, centered on the transit on 2 October 2019, during which we collected 12 spectra: 5 in-transit and 7 out-of-transit. We set the exposure time to 520 s, achieving a median S/N of 310.
TrES-1 b: We scheduled an observational campaign centered on the transits on 7 and 13 July 2020. In total, we collected 33 spectra: 9 in-transit and 24 out-of-transit; complemented with 9 P4 spectra. We set the exposure time to 820 s, achieving a median S/N of 97.
WASP-11 b/HAT-P-10 b: We collected 18 spectra on visit 1 (7 December 2019): 5 in-transit and 13 out-of-transit; and 12 spectra on visit 2 (18 December 2019): 7 in-transit and 5 out-of-transit. We also collected 13 P4 spectra. We set the exposure time to 820 s for both visits, achieving a median S/N of 105.
WASP-52 b: We collected 9 spectra on visit 1 (13 August 2019): 3 in-transit and 6 out-of-transit; and 12 spectra on visit 2 (10 October 2019): 5 in-transit and 7 out-of-transit. We also collected 13 P4 spectra. We set the exposure time to 820 s for visit 1 and 760 s for visit 2, achieving a median S/N of 85.
@@
WASP-69 b: We collected one useful in-transit spectrum on 20 September 2019, and 23 spectra on visit 2 (21 October 2019): 7 in-transit and 16 out-of-transit. We set the exposure time between 250–500 s, achieving a median S/N of 215.
WASP-76 b: We collected 22 spectra on visit 1 (2 October 2020): 10 in-transit and 12 out-of-transit; and 18 spectra on visit 2 (4 October 2020): 10 in-transit and 8 out-of-transit. We set the exposure time to 320 s for both visits, achieving a median S/N of 135.
WASP-80 b: We scheduled one visit, centered on the transit on 13 July 2019, during which we collected 11 spectra: 8 in-transit and 3 out-of-transit; complemented with 16 P4 spectra. We set the exposure time to 310 s, achieving a median S/N of 125.
WASP-127 b: We scheduled one visit, centered on the transit on 13 March 2019, during which we collected 35 spectra: 10 in-transit and 25 out-of-transit; complemented with 3 P4 spectra. We set the exposure time to 415 s, achieving a median S/N of 123.
WASP-156 b: We collected 14 spectra on visit 1 (24 October 2020): 6 in-transit and 8 out-of-transit; and 18 spectra on visit 2 (20 November 2020): 5 in-transit and 13 out-of-transit. We set the exposure time to 590 s for visit 2 and 550 s for visit 2, achieving a median S/N of 70 and 100, respectively.
WASP-177 b: We scheduled an observational campaign centered on the transits on 20 and 23 August 2020. In total, we collected 35 spectra: 8 in-transit and 27 out-of-transit. We set the exposure time to 830 s, achieving a median S/N of 85.
2.2 Telluric correction
@@
The main objective of these observations was the analysis of the planetary He i triplet line, which is surrounded by H2O absorption and OH emission lines from the Earth’s atmosphere. The relative positions of the He i triplet and telluric spectral lines change with the epoch due to the Doppler shift induced by the barycentric Earth radial velocity. Thus, we scheduled the campaigns at epochs that minimize the overlap of the He i planetary trace and the telluric lines (J. Orell-Miquel et al., 2022). Moreover, we also correct the science spectra from telluric lines to minimize their impact on the planetary signals. We use the python package muler (M. Gully-Santiago et al., 2022) to correct the target spectra from the telluric emission lines. The muler code uses the simultaneous sky spectra and accounts for the different efficiencies between target and sky fibers (Z. Zhang et al., 2023; M. Gully-Santiago et al., 2024). We use the telluric model from Goldilocks pipeline, which accounts for the instrumental profile and spectral resolution, to empirically fit and remove the telluric absorption lines (F. Yan et al., 2015; N. Casasayas-Barris et al., 2017).
2.3 Transmission spectrum analysis

### v2

@@ Abstract
The near-infrared helium triplet line is a powerful tool for studying atmospheric escape processes of close-in exoplanets, especially irradiated gas giants. Line profile fitting provides direct insight into the mechanisms driving atmospheric mass loss of close-in, Jupiter-sized planets. We present high-resolution transmission spectroscopy results for the helium triplet line of sixteen gas giants (). These observations are part of an extensive helium survey conducted using the Habitable Zone Planet Finder spectrograph on the 10 m Hobby-Eberly Telescope. For the first time, we provide constraints on the helium line for HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. Additionally, we are able to confirm previous robust or tentative detections for HD189733 b, HD209458 b, WASP-52 b, WASP-69 b, and WASP-76 b, and non-detections for HAT-P-3 b, WASP-11 b, WASP-80 b, WASP-127 b, and WASP-177 b. We do not confirm the previous helium narrow-band detection in HAT-P-26 b using high-resolution observations. To identify trends within the population of warm, hot, and ultra-hot Jupiters, we combined our results with available helium studies from the literature. As predicted by theory, we find that warm Jupiters with helium detections orbit K-type stars. However, the helium detections at equilibrium temperatures of 2000K are found in low-density planets orbiting F-type stars. We compiled a list of 46 irradiated gas giants, but more helium studies are needed to increase the sample and improve our understanding of atmospheric mass loss through helium observations.
@@ I Introduction
Gas giant exoplanets, particularly those on short‑period orbits, are the primary laboratories for atmospheric characterization because their large radii and extended atmospheres produce deep and readily detectable transit signals (Seager and Deming, 2010). Due to their proximity to the host star, they can reach high equilibrium temperatures (), ranging from warm Jupiters (WJs, 1000 K) through hot Jupiters (HJs, 1000–2000 K) to ultra‑hot Jupiters (UHJs, 2150 K, Parmentier et al., 2018). During a transit, starlight filters through the planetary atmosphere, enabling transmission spectroscopy with high-resolution spectrographs, which probes the composition of the upper atmospheric layers. The combination of large scale heights, high irradiation, and short orbital periods makes warm, hot, and ultra-hot Jupiters particularly favorable for atmospheric studies (Crossfield, 2015; Madhusudhan, 2019). These planets have provided the first detections of exoplanetary atomic species, and escaping material (Charbonneau et al., 2002; Vidal-Madjar et al., 2003).
@@
The atmospheric escape of close-in planets is primarily hydrodynamic, driven by the thermal energy deposited into the exoplanet’s atmosphere (Pierrehumbert, 2010; Koskinen et al., 2014). Photo-evaporation is the typical hydrodynamic escape process and is driven by stellar irradiation in the most energetic wavelengths, i.e., the X-ray and extreme ultraviolet (XUV111We define XUV as wavelengths up to 504 Å since this range is relevant to helium studies due to the ionization wavelength cutoff of the He i. We note that other works may define the XUV range different, e.g. up to the H ionization cutoff at 912Å.) range (Owen and Jackson, 2012). This energy input can inflate the upper atmosphere carrying light elements, such as hydrogen and helium, to higher altitudes. Additionally, the outer layers of the atmosphere of extremely close-in planets can extend beyond the Roche lobe, resulting in a high atmospheric mass loss rate (Owen, 2019). Roche lobe overflow escape is independent of stellar flux and can act alongside other evaporation mechanisms (Lecavelier des Etangs et al., 2004). In extreme cases, these extended, escaping atmospheres can result in the formation of comet‑like tails that interact with the stellar wind, detectable via atmospheric escape tracers, such as the He i triplet line (e.g., Zhang et al., 2023; Gully-Santiago et al., 2024). Over time, these mass‑loss processes can sculpt planetary demographics, contributing to observed features such as the sub‑Jovian desert, and providing empirical constraints for models of photo-evaporation (Owen and Jackson, 2012; Owen and Lai, 2018).
@@
The He i triplet at 10833 Å has become a key diagnostic for probing the upper atmospheres and escape processes of close‑in exoplanets, including irradiated gas giants (Seager and Sasselov, 2000; Oklopčić and Hirata, 2018; Lampón et al., 2021). The absorption arises from the metastable 23S state of neutral helium, which is populated when stellar XUV radiation ionizes helium atoms, and subsequent recombination fills the triplet state. However, the helium metastable state has a lifetime of 2.2 h (Drake, 1971). Thus, a constant and intense irradiation is needed to sustain detectability of the He i triplet line (Sanz-Forcada and Dupree, 2008). Since its first detection in WJs (Spake et al., 2018, WASP-107 b; Nortmann et al., 2018, WASP-69 b; and Allart et al., 2018, HAT-P-11 b), helium absorption has been reported in various types of exoplanets, providing a direct tracer of atmospheric escape accessible from ground-based telescopes. Moreover, ground-based high-resolution spectroscopy observations allow to retrieve several physical parameters from line profile fitting to study the hydrodynamical escape (e.g. Lampón et al., 2023).
Helium observations provide a unique window into both the instantaneous escape dynamics and the long‑term sculpting of exoplanet populations. This spectral line has been extensively studied in a wide range of exoplanets using narrow-band photometry (e.g., Vissapragada et al., 2022) and almost all ground-based near-infrared spectrographs (e.g., HPF, Ninan et al., 2020; NIRSPEC, Kirk et al., 2020; CRIRES+, Czesla et al., 2024; CARMENES, Nortmann et al., 2018; Allart et al., 2018; GIANO-B, Guilluy et al., 2023; SPIRou, Masson et al., 2024; Allart et al., 2023; NIRPS, Allart et al., 2025a; and IRD, Kawauchi et al., 2022). Moreover, the convenience of observing the He line has been leveraged for large-scale surveys aimed at understanding the atmospheric retention, evolution, and population trends of exoplanets (e.g. Vissapragada et al., 2022; Guilluy et al., 2023; Allart et al., 2023; Orell-Miquel et al., 2024).
@@
In this context, this work presents the results of an observational study aimed at characterizing atmospheric escape in gas giants. We have been carrying out a survey targeting the He i triplet line for planets with 0.5 RJup and ranging 800–2100 K around F-, G-, and K-type host stars. Figure 1 presents the planets observed by our survey. The individual analyses and discussions for HAT-P-32 b and HAT-P-67 b are presented in Zhang et al. (2023), and Gully-Santiago et al. (2024), respectively. Here, we only included those results, and not new analysis on these two systems in our helium detections discussion.
II Observations and data analysis
II.1 HPF spectroscopic observations
@@
We obtained spectroscopic observations of gas giant host stars with the Habitable Zone Planet Finder (HPF; Mahadevan et al. 2012, 2014; Metcalf et al. 2019) mounted on the 10 m Hobby-Eberly Telescope (HET; Ramsey et al. 1998; Shetrone et al. 2007). HPF spectrograph covers the near-infrared wavelength range from 0.81–1.38 m with a resolving power of 55 000. HPF data were reduced using the Goldilocks222https://github.com/grzeimann/Goldilocks_Documentation pipeline (Ninan et al., 2018; Kaplan et al., 2019), obtaining the science, sky, and calibration. To avoid saturation or scattered light from the laser frequency comb, we did not use a simultaneous calibration source in the calibration fiber.
@@
The HET telescope structure can rotate in azimuth but it is fixed at an elevation of 55 degrees. Target tracking is performed by a prime focus instrument package that moves across the focal plane to follow the target. Due to the design of the HET, we are limited to observations during “tracks” of typically 1 hour before (east track) and after (west track) target culmination, making it difficult observe a full transit, or to obtain in-transit and out-of-transit spectra on the same night. We organized HPF observations into visit campaigns around an exoplanet transit. We obtained in-transit spectra (IN) along with out-of-transit spectra (OUT) from surrounding nights. We also obtained random-phase spectra, referred to as P4 observations, to monitor long-term stellar variability. P4 observations are typically used to fill the HET queue or when the conditions are not ideal for higher-priority targets. Although P4 spectra are out-of-transit, they are not considered OUT because they are collected away from IN visits.
@@
Table 3 provides the log of the observations for our targets, with the list of observed tracks, exposure time, and signal-to-noise ratio (S/N) per extracted pixel of the échelle order where the He triplet line is located. We provide a summary of the observations performed for each target:
HAT-P-3 b: We scheduled one visit with 16 spectra: 8 in-transit (1 February 2021) and 8 out-of-transit spectra. We additionally obtained 4 P4 spectra.
HAT-P-12 b: We collected 15 spectra on visit 1: 5 in-transit (31 March 2020) and 10 out-of-transit; and 12 spectra on visit 2: 8 in-transit (27 May 2025) and 4 out-of-transit. We also collected 2 P4 spectra.
HAT-P-17 b: We collected 23 spectra on visit 1: 11 in-transit (8 June 2020) and 12 out-of-transit; and 32 spectra on visit 2: 10 in-transit (30 July 2020) and 22 out-of-transit. We also collected 8 P4 spectra.
HAT-P-26 b: We observed one visit with 13 spectra: 5 in-transit (1 February 2021) and 8 out-of-transit spectra. We additionally obtained 2 P4 spectra.
@@
HD118203 b: We collected 26 spectra on visit 1: 18 in-transit (7 January 2021) and 8 out-of-transit; and 29 spectra on visit 2: 18 in-transit (19 February 2021) and 11 out-of-transit. We additionally obtained 2 P4 spectra.
HD189733 b: We obtained 9, 9, and 2 in-transit spectra for visits 1 (4 May 2019), 2 (13 June 2019), and 3 (1 August 2020), respectively. We also collected 11 spectra as out-of-transit in visit 4, complemented with 10 P4 spectra (two of them are later used as out-of-transit for visit 1).
HD209458 b: We collected 6 spectra on visit 1: 5 in-transit (2 october 2019), and 1 out-of-transit; and 6 out-of-transit spectra on visit 2.
TrES-1 b: We obtained in total 33 spectra: 5 in-transit spectra on visit 1 (7 July 2020), 5 in-transit spectra on visit 2 (13 July 2020), and 23 out-of-transit spectra contemporaneous to both visits. We additionally obtained 9 P4 spectra.
WASP-11 b/HAT-P-10 b: We collected 18 spectra on visit 1: 5 in-transit (7 December 2019) and 13 out-of-transit; and 12 spectra on visit 2: 5 in-transit (18 December 2019) and 5 out-of-transit. We also collected 8 P4 spectra.
WASP-52 b: We collected 9 spectra on visit 1: 4 in-transit (13 August 2019) and 5 out-of-transit; and 12 spectra on visit 2: 5 in-transit (10 October 2019) and 7 out-of-transit.
@@
WASP-69 b: We collected a single in-transit spectrum on visit 1 (20 September 2019), and 23 spectra on visit 2: 7 in-transit (21 October 2019) and 16 out-of-transit.
WASP-76 b: We collected 22 spectra on visit 1: 10 in-transit (2 October 2020) and 12 out-of-transit; and 18 spectra on visit 2: 10 in-transit (4 October 2020) and 8 out-of-transit.
WASP-80 b: We observed one visit with 11 spectra: 8 in-transit (13 July 2019) and 3 out-of-transit spectra. We additionally obtained 16 P4 spectra.
WASP-127 b: We observed one visit with 35 spectra: 10 in-transit (13 March 2019) and 25 out-of-transit spectra. We additionally obtained 4 P4 spectra.
WASP-156 b: We collected 14 spectra on visit 1: 6 in-transit (24 October 2020) and 8 out-of-transit; and 18 spectra on visit 2: 5 in-transit (20 November 2020) and 13 out-of-transit.
WASP-177 b: We obtained in total 35 spectra: 4 in-transit spectra on visit 1 (20 August 2020), 4 in-transit spectra on visit 2 (23 August 2020), and 27 out-of-transit spectra contemporaneous to both visits.
II.2 Telluric correction
@@
The main objective of these observations was the analysis of the planetary He i triplet line, which is surrounded by H2O absorption and OH emission lines from the Earth’s atmosphere. The relative positions of the He i triplet and telluric spectral lines change with the epoch due to the Doppler shift induced by the barycentric Earth radial velocity. Thus, we mainly scheduled the campaigns at epochs that minimize the overlap of the He i planetary trace and the telluric lines (Orell-Miquel et al., 2022). Moreover, we also correct the science spectra from telluric lines to minimize their impact on the planetary signals. We use the python package muler (Gully-Santiago et al., 2022) to correct the target spectra from the telluric emission lines. The muler code uses the simultaneous sky spectra and accounts for the different efficiencies between target and sky fibers (Zhang et al., 2023; Gully-Santiago et al., 2024). We remove the telluric absorption lines using a telluric model computed from HPF/HET observations333The model is available at the Goldilocks pipeline documentation (telluricpca.fits)., thus accounts for the instrumental profile and spectral resolution. For each individual spectra, we empirically fit the strongest H2O lines to account for strength line differences between the model and the data. The observed spectrum is divided by the scaled model, removing the telluric absorption features (e.g., Yan et al., 2015; Casasayas-Barris et al., 2017).
II.3 Transmission spectrum analysis